OpenStudy (anonymous):

solve for x. logx + log(x-3)=1

OpenStudy (anonymous):

you can rewrite it as log of product

OpenStudy (anonymous):

The sum of logs is the log of the product. \[log(a) + log(b) = log(a\cdot b)\]

OpenStudy (anonymous):

So you can use that to re-write your equation as: \[log(x(x-3)) = 1\]

OpenStudy (anonymous):

Now try raising 10 to the power of both sides.

OpenStudy (anonymous):

great thanks a ton